60 mi/2 h = 180 mi/?

Found 2 solutions by stanbon, ReadingBoosters:
Answer by stanbon(75887)

60 mi/2 h = 180 mi/?
-----
Cross multiply to get:
60*? = 2*180
? = 2*3
? = 6 hours
============
Cheers,
Stan H.

Answer by ReadingBoosters(3246)

60x = 2(180)
60x = 360
x = 6 hours
.....................
Delighted to help; feel free to email if you have further questions.
HomeworkHelpers@readingboosters.com
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.